OpenStudy (anonymous):

Part 1: Explain, in complete sentences, whether or not the expression x2 + 24x + 144 fits one of the special patterns. If so, which one? (2 points) Part 2: Explain how the polynomial is factored. (2 points) Part 3: Provide the factors. (2 points)

jimthompson5910 (jim_thompson5910):

were you able to get anywhere with this one?

OpenStudy (anonymous):

nope-_-

jimthompson5910 (jim_thompson5910):

ok let's try to factor x^2 + 24x + 144

jimthompson5910 (jim_thompson5910):

can you find me two numbers that multiply to 144 (last term) AND add to 24 (middle coefficient)?

jimthompson5910 (jim_thompson5910):

are you able to find two such numbers?

OpenStudy (anonymous):

im horrible at this stufff..... helppp

jimthompson5910 (jim_thompson5910):

ok how about 1 and 144 1*144 = 144 ...yes 1+144 = 145 ... no (we want them to add to 24)

jimthompson5910 (jim_thompson5910):

2 and 72? 2*72 = 144 .... yes 2 + 72 = 74 ... no

jimthompson5910 (jim_thompson5910):

3 and 48? 3*48 = 144 ... yes 3+48 = 51 ... no

jimthompson5910 (jim_thompson5910):

you keep going until you have the two numbers add to 24

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yeahh.. im still lost.

jimthompson5910 (jim_thompson5910):

ok it turns out 12 and 12 both multiply to 144 and add to 24

jimthompson5910 (jim_thompson5910):

so x^2 + 24x + 144 factors to (x+12)(x+12)

jimthompson5910 (jim_thompson5910):

which is basically a perfect square since (x+12)(x+12) is the same as (x+12)^2

OpenStudy (anonymous):

alright

jimthompson5910 (jim_thompson5910):

a useful formula is this (ax+b)^2 = a^2*x^2 + 2*a*b*x + b^2 which is a formula used to expand out any perfect square

jimthompson5910 (jim_thompson5910):

and it turns out that x^2 + 24x + 144 is in that form above

jimthompson5910 (jim_thompson5910):

does that help? or are you still stuck?

OpenStudy (anonymous):

Alright..

OpenStudy (anonymous):

& yeah I kinda am.

jimthompson5910 (jim_thompson5910):

ok where are you stuck

OpenStudy (anonymous):

the wholee thingg.

jimthompson5910 (jim_thompson5910):

you saw how I factored x^2 + 24x + 144 into (x+12)^2 right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so that shows us that x^2 + 24x + 144 is a perfect square trinomial

jimthompson5910 (jim_thompson5910):

all perfect square trinomials are of the form (ax+b)^2 = a^2*x^2 + 2*a*b*x + b^2

jimthompson5910 (jim_thompson5910):

in this particular case, a = 1 and b = 12

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so we can see that if we plug in a = 1 and b = 12, we get (ax+b)^2 = a^2*x^2 + 2*a*b*x + b^2 (1*x+12)^2 = (1)^2*x^2 + 2*1*12*x + (12)^2 (x+12)^2 = x^2 + 24x + 144

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so that's what they mean by special patterns

jimthompson5910 (jim_thompson5910):

basically it fits the form a^2*x^2 + 2*a*b*x + b^2, which is a perfect square

jimthompson5910 (jim_thompson5910):

starting to get clearer (a bit)?

OpenStudy (anonymous):

yeah a little

jimthompson5910 (jim_thompson5910):

ok just keep practicing and it'll become even more clearer (hopefully)

jimthompson5910 (jim_thompson5910):

so let's answer this one by one (part by part)

jimthompson5910 (jim_thompson5910):

Part 1: Explain, in complete sentences, whether or not the expression x2 + 24x + 144 fits one of the special patterns. If so, which one? (2 points) Part 2: Explain how the polynomial is factored. (2 points) Part 3: Provide the factors. (2 points)

jimthompson5910 (jim_thompson5910):

Part 1: Explain, in complete sentences, whether or not the expression x2 + 24x + 144 fits one of the special patterns. If so, which one? (2 points) it fits the form a^2*x^2 + 2*a*b*x + b^2 where a = 1 and b = 12 this basic form is a perfect square trinomial

OpenStudy (anonymous):

alright

jimthompson5910 (jim_thompson5910):

Part 2: Explain how the polynomial is factored. (2 points) I'll let you answer this part. Use the formula I provided to get you started

jimthompson5910 (jim_thompson5910):

able to get anywhere with it?

OpenStudy (skullpatrol):

@jim_thompson5910 nice job :)